OpenStudy (anonymous):

Find a formula for the described function. A rectangle has perimeter 16 m. Express the area A of the rectangle as a function of the length, L, of one of its sides.

OpenStudy (anonymous):

Is there any more information? Do they give you any numbers for the sides of the rectangle?

OpenStudy (anonymous):

no, its just that unfortunately

OpenStudy (loser66):

perimeter of a rectangular is defined by formula (L + W )/2 and it is = 16. That means \[\frac{L+W}{2}=16\\L+W=32 \\W = 32-L\] Now, The area of rectangular is defined by A= L*W , replace W = 32 -L into this, you have the function of area of A respect to L is A = L *(32 -L) or A = 32L-L^2

OpenStudy (anonymous):

it's not 32, it's 8.

OpenStudy (loser66):

oh ya, you are right, my bad, (L+W)*2 not /2 I am sorry.